# Generalized Uncertainty Principle theories and their classical interpretation

### v1

@@ Abstract
Finally, we demonstrate that a properly defined GUP theory can be correctly interpreted classically if and only if the corresponding quantum commutators satisfy the Jacobi identities.
@@ 1 Introduction
Several arguments, stemming from String Theory and Gedanken experiments [1, 2, 3], suggest that a modification of the HUP is indeed necessary at some level. It is straightforward to realize that the most rigorous way to implement such a modification is to properly induce a different structure of the Heisenberg algebra, i.e. alter the canonical commutation relations between quantum conjugate operators. According to the type of modification introduced in the algebra, different classes of GUP theories can be identified. In the class of interest in this paper, whose structure will be discussed below, the main physical consequences due to the constructed deformation are twofold: first, the appearance of a minimum structure in the configuration space of the theory, e.g., a minimum length in position space, in the form of a non-zero minimal uncertainty in the configuration operators; second, a resulting non-commutativity between the configuration operators, which can be interpreted as the emergence of a non-commutative “geometry” in configuration space [4, 5, 6, 7].
@@
In this respect, even if applications of GUP theories can be found in countless physical scenarios [8], perhaps the most relevant implementation of these effective models concerns early-stage cosmology [9, 10, 11] and black hole physics [12]. In these contexts, they can introduce a drastically different dynamics at the Planck scale, providing useful physical insights on the role played by quantum gravity.
@@
In light of this, in this paper, we investigate if it is possible to correctly implement a GUP theory at the classical level, using the symplectic geometry formulation. Given a even-dimensional smooth manifold $\mathcal{M}$, we interpret the variables $(q,p)$ as coordinates on $\mathcal{M}$ and define a 2-form which induces the desired GUP Poisson brackets, requiring this 2-form to be symplectic, i.e., non-degenerate and closed. This allows us to completely characterize the symplectic structure of the theory in order to be consistent, finding precise and completely general relations between the phase-space functions which control the deformation of the Poisson structure. Consistently, the results we obtain are entirely equivalent to imposing the validity of Jacobi identities for the deformed Poisson brackets, exactly as it should be.
On the basis of these general symplectic relations, we further constrain the Poisson brackets between the generalized coordinates. Specifically we require that they form a set of phase-space functions, whose algebra corresponds to that of an angular momentum. This request, commonly imposed in many GUP theories, helps to maintain the rotational invariance of the algebra. This allows us to further specify the structure of the Poisson brackets and determine the generator of rotations in these theories. Our results are essentially compatible with those of [17, 18], where a similar analysis is conducted at the quantum level using the Jacobi identities. However, while [17, 18] suggest that the angular momentum operator constrains the theory additionally due to the presence of the spin, our analysis shows that this possibility is ruled out at the classical level. Specifically, we demonstrate that an angular momentum, which proper generates the rotations, exists as a phase-space function but, since no analogue of the spin emerges classically, does not impose any other constraint.
@@
The complete general determination of the symplectic structure of the classical theory ultimately allows us to demonstrate an important result that concludes this work: a GUP theory is consistently implemented at the classical level if and only if the corresponding quantum commutators satisfy the Jacobi identities.
@@
Conversely, if the classical relations that ensure the theory is symplectic are satisfied, meaning the classical Jacobi identities hold, it can be demonstrated that the corresponding quantum commutators will also satisfy these identities.
As we will see, this second point is delicate due to the operator ordering ambiguities that arise in the quantization of certain phase space functions. We will show that some of the most natural orderings satisfy the Jacobi identities, thereby providing a guideline for the prescription of the theory at the quantum level.
The paper is organized as follows: Section 2 reviews fundamental elements and concepts of symplectic geometry relevant to the formulation of Hamiltonian classical mechanics. Moreover, we introduce the concept of classical interpretation of a GUP theory and the main Theorem associated with this definition.
In Section 3, we derive the symplectic structure of a general GUP theory within the class of interest, based on its Poisson brackets. We particularly focus on the relationships dictated by the closed condition of the 2-form. We perform computations in dimensions ranging from 1d to d-dimensional cases, providing concrete examples of GUP theories that align perfectly with the derived relations, including, e.g. the GUP model found in [4], along with explicit calculations.
Section 4 we consider the case where an angular momentum function defines the Poisson brackets between generalized coordinates and we suitably compare our findings with those presented in [17, 18]. In Section 5 we give a demonstration of the Theorem presented in Section 2 which clarify the relation between the classical and quantum framework of these theories. Finally, in Section 6 we give our conclusions, emphasizing the significance of clarifying the consistency of GUP theories at the classical level for any semiclassical applications.
@@ 2 Classical interpretation and Jacobi identity
A symplectic form $\omega$ induces a Lie algebra structure on $\mathcal{C}^{\infty}(\mathcal{M})$ via the Poisson brackets. Let $f,g\in\mathcal{C}^{\infty}(\mathcal{M})$, and let $X_{f}$ be Hamiltonian vector field of $f$ defined by $df=X_{f}\mathbin{\lrcorner}\omega$, the Poisson bracket of two functions are $\{f,g\}=-\omega(X_{f},X_{g})$. The Poisson brackets satisfy the Jacobi identity and define a nondegenerate Poisson structure on $\mathcal{M}$ [16].
@@
The inverse of that matrix is immediate and, so, we obtain the matrix of the symplectic form
@@
However, the 2-form in (2.4) is not a symplectic form a priori. This classical structure is a relevant feature for the study of quantum physical theory. Indeed, despite in the literature it is debated if the semiclassical dynamics of a GUP is implemented by its classical interpretation, we can anyway relate the algebraic properties of the quantum commutation relations to the symplectic structure. When $L_{ij}$ and $f$ can be interpreted as smooth functions of the coordinates on an open subset $\mathcal{U}\subset\mathbb{R}^{2d}$, and $f$ is strictly positive on $\mathcal{U}$, the following Theorem holds:
Theorem 2.3.
@@
The proof of this theorem is given in Section 5. As we will see, the proof provide us with a correct prescription for the quantization of some relevant phase-space functions, selecting suitable operator orderings among all the possible ones.
The following Section shows how to construct the symplectic form $\omega$ just imposing the non-degeneracy and the closure of $\omega$, funding some constraints for the expressions of $L$ and $f$. Because of Theorem 2.3, this construction is valid also at the quantum level. The considerations expressed in Section 3 will constitute the main part of the proof of Theorem 2.3.
@@ 3 Building the function f(p)
The $1d$ case is highly trivial. Being, in this case, the phase space a 2-dimensional manifold, every 2-form is closed. This means that the closed condition is automatically satisfied. From this, immediately follows that every 1-dimensional GUP theory, with a suitable $f$ respecting the minimal conditions imposed above, has a symplectic structure and hence a classical interpretation.
@@
These relations can be read as a gradient equation for $f$ in an open set $U\subset\mathbb{R}^{2}$ with respect to $p_{1},p_{2}$ variables. Together with the independence of $f$ from $q_{1},q_{2}$, from (3.3) we can deduce a general form for $l$ function:
@@
3.2.2 Interesting example: Kempf-Mangan-Mann 2D-GUP
@@
If we now fix $k=-2\beta$, where $\beta$ is a positive parameter, what we obtain is the well-known GUP model first proposed by Kempf, Mangano, and Mann in [4], at the quantum level.
@@ 3.2.3 Explicit computation: Polynomial functions
We can check a posteriori that $f$ is smooth and, if $c>0$, there exist $\varepsilon>0$ such that $f$ is strictly positive in $\mathbb{B}_{\varepsilon}^{2}\subset\mathbb{R}^{2}$. Thus, our phase space is $\mathcal{M}=\mathbb{R}^{2}\times\mathbb{B}_{\varepsilon}^{2}$ with a symplectic form given in coordinates by (3.1), where the functions $l$ and $f$ are defined in (3.2.3) and (3.11), respectively.
@@
$\tilde{\omega}_{ij}=\begin{pmatrix}0&-\mathrm{id}_{d}\\ \mathrm{id}_{d}&hL\end{pmatrix}.$
@@
That is the independence of $f$ on $q$s. We can show it by noticing that the equations are antisymmetric in $i,j$ and the only non-trivial equations are for $i=k$ (equivalently $j=k$). providing a set of $d$ independent equations
@@
Hence, considering the gradient equation satisfied, the strange equation becomes a constraint on the $S_{ij}$ functions, which are not fixed by the form of $f(p)$
@@ 3.3.1 The 3-dimensional case
Here, $S_{ij},P,Q,R$ are smooth functions of $p_{1},p_{2},p_{3}$. In such a way, the system (3.19) is recast as 3 independent equations
@@
The functions $S_{ij}$ are solution of the above equation if they are solution of the following system
@@
Hence, the solution is $S_{ij}=\pm f^{2}e^{c_{ij}(p_{i},p_{j})}$, where $c_{ij}=c_{ji}$ are arbitrary functions on $p_{i},p_{j}$ only.
@@ 3.3.2 Interesting example: Kempf-Mangano-Mann 3D-GUP
The Kempf-Mangano-Mann (KMM) GUP theory is recovered for $c=1$ and $S_{ij}\equiv 0$. However, as we already said, this is not the unique choice for $S_{ij}$. Nevertheless, considering $\beta$ the control parameter and requiring that the usual Heisenberg algebra is recovered for $\beta\to 0$, it is possible to fix the $S_{ij}$ functions. Indeed, $S_{ij}=f^{2}e^{c_{ij}(p_{i},p_{j})}$ is an admissible solution of 3.18 but it does not have the correct limit because $f\xrightarrow[\beta\to 0]{}c$, forcing $c=1$, and so $S_{ij}\xrightarrow[\beta\to 0]{}e^{c_{ij}(p_{i},p_{j})}$, and it cannot be zero.
@@ 4 Maggiore GUP as an application
The mathematical relations we have derived in Section 3 are completely general and are based only on the request the 2-form $\omega$ (2.4) is symplectic, without further assumptions on the $f$ and $L_{ij}$ functions.
Nevertheless, usually, in constructing a GUP theory other constrains are imposed on the general algebra (2). The most common additional demands include the rotational invariance of the algebra and the condition that the $L_{ij}$ functions satisfy an angular momentum algebra.
@@
Since the solution with the minus sign violates the condition of having an $f$ function with definite sign, as we establish at the beginning of the paper, in order to deal with a symplectic form defined everywhere, in the following we are going to consider only the solution with the plus sign, obtained for $a(\hat{p})=-1$.
@@
$[\hat{q}_{i},\hat{q}_{j}]=-\frac{\hbar^{2}}{\kappa^{2}}i\epsilon_{ijk}\hat{J}_{k},$
$[\hat{q}_{i},\hat{p}_{j}]=i\hbar\delta_{ij}\sqrt{1+\tfrac{\hat{\rho}^{2}}{\kappa^{2}c^{2}}}.$
As it should be clear from the derivation, this specific GUP model acquires relevance since it is one of the two most general models, defined only by the imposition of Jacobi identities and the inclusion of the spin operator. We stress that the generality of the model has to be intended in the sense specified above.
@@ 4.1 Closure condition and Maggiore algebra
$\begin{cases}(145)&\frac{\partial f}{\partial p_{2}}+a(p)\frac{\partial J_{3}}{\partial q_{1}}=0,\\ (146)&\frac{\partial f}{\partial p_{3}}-a(p)\frac{\partial J_{2}}{\partial q_{1}}=0,\\ (156)&\frac{\partial J_{1}}{\partial q_{1}}=0,\\ (245)&\frac{\partial f}{\partial p_{1}}-a(p)\frac{\partial J_{3}}{\partial q_{2}}=0,\\ (246)&\frac{\partial J_{2}}{\partial q_{2}}=0,\\ (256)&\frac{\partial f}{\partial p_{3}}+a(p)\frac{\partial J_{1}}{\partial q_{2}}=0,\\ (345)&\frac{\partial J_{3}}{\partial q_{3}}=0,\\ (346)&\frac{\partial f}{\partial p_{1}}+a(p)\frac{\partial J2}{\partial q_{3}}=0,\\ (356)&\frac{\partial f}{\partial p_{2}}-a(p)\frac{\partial J_{1}}{\partial q_{3}}=0.\end{cases}$ (4.5)
@@
Accordingly, the gradient equation for $f$ that can be extracted from the system (4.5), becomes:
@@
By means of the expressions (4.1) for the $J$s functions we obtain a system of equations, which compactly we can write as:
@@
As a homogeneous linear system, its determinant is $f^{-2}(f^{2}+a(p_{1}^{2}+p_{2}^{2}+p_{3}^{2}))$. The singular points are represented by the equation
$a(p)(p_{1}^{2}+p_{2}^{2}+p_{3}^{2})=-f^{2}(p),$
which describes a closed surface in $\mathbb{R}^{3}$. Clearly, this equation admits solution only for $a(p)<0$. Where in $\mathbb{R}^{3}$ the determinant is not singular, the unique solution of (4.14) is $\mathbf{s}=0$. Hence, since $s_{i}$ is a smooth function everywhere null but on a closed surface, for continuity it must be zero everywhere in $\mathbb{R}^{3}$.
@@
By employing the last obtained form of the $J^{\prime}s$ we get:

### v2

@@ Abstract
Finally, we postulate that a properly defined GUP theory can be correctly interpreted classically if and only if the corresponding quantum commutators satisfy the Jacobi identities, identifying those quantization prescriptions for which this holds true.
@@ 1 Introduction
Several arguments, stemming from String Theory and Gedanken experiments [1, 2, 3], suggest that a modification of the HUP is indeed necessary at some level. It is straightforward to realize that the most rigorous way to implement such a modification is to properly induce a different structure of the Heisenberg algebra, i.e. alter the canonical commutation relations between quantum conjugate operators. According to the type of modification introduced in the algebra, different classes of GUP theories can be identified. In the class of interest in this paper, whose structure will be discussed below, the main possible physical consequences due to the constructed deformation are twofold: first, the appearance of a minimum structure in the configuration space of the theory, e.g. a minimum length in position space, in the form of a non-zero minimal uncertainty in the configuration operators; second, a resulting non-commutativity between the configuration operators, which can be interpreted as the emergence of a non-commutative “geometry” in configuration space [4, 5, 6, 7].
@@
In this respect, even if applications of GUP theories can be found in countless physical scenarios [8], perhaps the most relevant implementation of these effective models concerns early-stage cosmology [9, 10, 11] and black hole physics [12, 13]. In these contexts, they can introduce a drastically different dynamics at the Planck scale, providing useful physical insights on the role played by quantum gravity.
@@
In light of this, in this paper, we investigate if it is possible to correctly implement a GUP theory at the classical level, using the symplectic geometry formulation. Given an even-dimensional smooth manifold $\mathcal{M}$, we interpret the variables $(q,p)$ as coordinates on $\mathcal{M}$ and define a 2-form which induces the desired GUP Poisson brackets, requiring this 2-form to be symplectic, i.e., non-degenerate and closed. This allows us to completely characterize the symplectic structure of the theory in order to be consistent, finding precise and completely general relations between the phase-space functions which control the deformation of the Poisson structure. Consistently, the results we obtain are entirely equivalent to imposing the validity of Jacobi identities for the deformed Poisson brackets, exactly as it should be.
On the basis of these general symplectic relations, we further constrain the Poisson brackets between the generalized coordinates. Specifically we require that they form a set of phase-space functions, whose algebra corresponds to that of an angular momentum. This request, commonly imposed in many GUP theories, helps to maintain the rotational invariance of the algebra. This allows us to further specify the structure of the Poisson brackets and determine the generator of rotations in these theories. Our results are essentially compatible with those of [18, 19], where a similar analysis is conducted at the quantum level using the Jacobi identities. However, while [18, 19] suggests that the angular momentum operator constrains the theory additionally due to the presence of the spin, our analysis shows that this possibility is ruled out at the classical level. Specifically, we demonstrate that an angular momentum, which proper generates the rotations, exists as a phase-space function but, since no analogue of the spin emerges classically, does not impose any other constraint.
@@
The complete general determination of the symplectic structure of the classical theory ultimately allows us to formulate a Conjecture that concludes this work: a GUP theory is consistently implemented at the classical level if and only if the corresponding quantum commutators satisfy the Jacobi identities.
@@
Conversely, if the classical relations that ensure the theory is symplectic are satisfied, meaning that the classical Jacobi identities hold, it is not obvious that the quantum Jacobi identities are satisfied due to the ambiguities in the quantization procedures. Nevertheless, we show that some of the most natural ordering prescriptions satisfy the Jacobi identities, thereby providing a potential guideline for the prescription of the theory at the quantum level.
The paper is organized as follows: Section 2 reviews fundamental elements and concepts of symplectic geometry relevant to the formulation of Hamiltonian classical mechanics. Moreover, we introduce the concept of classical interpretation of a GUP theory and the main Conjecture associated with this definition.
In Section 3, we derive the symplectic structure of a general GUP theory within the class of interest, based on its Poisson brackets. We particularly focus on the relationships dictated by the closure condition of the 2-form. We perform computations ranging from 1 to d-dimensional cases, providing concrete examples of GUP theories that align perfectly with the derived relations, including, e.g. the GUP model found in [4], along with explicit calculations.
In Section 4 we consider the case where an angular momentum function defines the Poisson brackets between generalized coordinates and we suitably compare our findings with those presented in [18, 19].
In Section 5 we present a wide class of cases for which the Conjecture presented in Section 2 holds. The Conjecture provides a relation between the classical and quantum framework of these theories.
Finally, in Section 6 we give our conclusions, emphasizing the significance of clarifying the consistency of GUP theories at the classical level for any semiclassical applications.
@@ 2 Classical interpretation and Jacobi identity
A symplectic form $\omega$ induces a Lie algebra structure on $\mathcal{C}^{\infty}(\mathcal{M})$ via the Poisson brackets. Let $f,g\in\mathcal{C}^{\infty}(\mathcal{M})$, and let $X_{f}$ be Hamiltonian vector field of $f$ defined by $df=X_{f}\mathbin{\lrcorner}\omega$, the Poisson brackets of two functions are $\{f,g\}=-\omega(X_{f},X_{g})$. The Poisson brackets satisfy the Jacobi identity and define a non-degenerate Poisson structure on $\mathcal{M}$ [17].
@@
The inverse of this matrix is immediate and so we obtain the matrix of the symplectic form
@@
However, the 2-form in (2.4) is not a symplectic form a priori. This classical structure is a relevant feature for the study of quantum physical theory. Indeed, despite in the literature it is debated if the semiclassical dynamics of a GUP is implemented by its classical interpretation, we can anyway relate the algebraic properties of the quantum commutation relations to the symplectic structure. When $L_{ij}$ and $f$ can be interpreted as smooth functions of the coordinates on an open subset $\mathcal{U}\subset\mathbb{R}^{2d}$, and $f$ is strictly positive on $\mathcal{U}$, we can formulate the following Conjecture:
Conjecture 2.3.
@@
In Section 5 we provide a series of cases in which the Conjecture is verified. As we will see, we can postulate that the Conjecture provides us with a correct prescription for the quantization of some relevant phase-space functions, selecting suitable operator orderings among all the possible ones.
The following Section shows how to construct the symplectic form $\omega$ just imposing the non-degeneracy and the closure of $\omega$, funding some constraints for the expressions of $L$ and $f$. The considerations expressed in Section 3 will constitute the main part of the construction of the Conjecture 2.3.
@@ 3 Building the function f(p)
The 1-dimensional case is highly trivial. Being, in this case, the phase space a 2-dimensional manifold, every 2-form is closed. This means that the closed condition is automatically satisfied. From this, immediately follows that every 1-dimensional GUP theory, with a suitable $f$ respecting the minimal conditions imposed above, has a symplectic structure and hence a classical interpretation.
@@
These relations can be read as a gradient equation for $f$ in an open set $U\subset\mathbb{R}^{2}$ with respect to $p_{1},p_{2}$ variables. Together with the independence of $f$ from $q_{1},q_{2}$, from (3.3) we can deduce a general form for the $l$ function:
@@
3.2.2 Interesting example: Kempf-Mangano-Mann 2D-GUP
@@
Clearly we are not considering a possible dependence of the integration constant on the control parameter $\kappa$. The idea is indeed to introduce the deformation parameter in a natural way only once and let it properly propagate along the equations.
If we now fix $\kappa=-2\beta$, where $\beta$ is a positive parameter, what we obtain is the well-known GUP model first proposed by Kempf, Mangano, and Mann in [4], at the quantum level.
@@
It is worth to notice that the symplectic form naturally induces a volume form in the phase space, which in the 2-dimensional case reads as $\mathrm{Vol}_{\omega}:=\omega\wedge\omega=2!f^{-2}\ dq^{1}\wedge dq^{2}\wedge dp^{1}\wedge dp^{2}$.
This volume form is clearly well-defined if and only if the $\omega$ is non-degenerate, that is if $f\neq 0$. The resulting condition is very close to the one obtained in the quantum framework from the request of symmetry of the fundamental operators, in particular the position operator. In this context in order to recover the symmetry of the operators - necessary but not sufficient condition for the self-adjointness - we need to deform the Lebesgue measure in a very similar way, introducing a factor $1/f$. As a consequence, both measures result to be well-defined if and only if $f\neq 0$, which in turn is a condition related, in the classical context, to the non-degeneracy of $\omega$.
@@ 3.2.3 Explicit computation: Polynomial functions
We can check a posteriori that $f$ is smooth and, if $c>0$, there exists $\varepsilon>0$ such that $f$ is strictly positive in $\mathbb{B}_{\varepsilon}^{2}\subset\mathbb{R}^{2}$. Thus, our phase space is $\mathcal{M}=\mathbb{R}^{2}\times\mathbb{B}_{\varepsilon}^{2}$ with a symplectic form given in coordinates by (3.1), where the functions $l$ and $f$ are defined in (3.2.3) and (3.11), respectively.
@@
$\tilde{\omega}_{ab}=\begin{pmatrix}0&-\mathrm{id}_{d}\\ \mathrm{id}_{d}&hL\end{pmatrix}.$
@@
That is the independence of $f$ on $q$s. We can show it by noticing that the equations are antisymmetric in $i,j$ and the only non-trivial equations are for $i=k$ (equivalently $j=k$), providing a set of $d$ independent equations
@@
Hence, considering the gradient equation satisfied, the strange equation becomes a constraint on the $S_{ij}$ functions, which are not fixed by the form of $f(p)$:
@@ 3.3.1 The 3-dimensional case
Here, $S_{ij},P,Q,R$ are smooth functions of $p_{1},p_{2},p_{3}$. In such a way, the system (3.19) is recast as three independent equations
@@
We note that solutions of the following system are solutions of the above equation as well:
@@
Hence, a possible solution is $S_{ij}=\pm f^{2}e^{c_{ij}(p_{i},p_{j})}$, where $c_{ij}=c_{ji}$ are arbitrary functions on $p_{i},p_{j}$ only.
@@ 3.3.2 Interesting example: Kempf-Mangano-Mann 3D-GUP
The Kempf-Mangano-Mann (KMM) GUP theory is recovered for $c=1$ and $S_{ij}\equiv 0$. However, as we already said, this is not the unique choice for $S_{ij}$. Nevertheless, considering $\beta$ the control parameter and requiring that the usual Heisenberg algebra is recovered for $\beta\to 0$, it is possible to exclude some $S_{ij}$ functions. For instance, $S_{ij}=f^{2}e^{c_{ij}(p_{i},p_{j})}$ is an admissible solution of 3.18 but it does not have the correct limit because $f\xrightarrow[\beta\to 0]{}c$, forcing $c=1$, and so $S_{ij}\xrightarrow[\beta\to 0]{}e^{c_{ij}(p_{i},p_{j})}$, which cannot be zero. For the same reason discussed in Section 3.2.2, here we are not considering a possible dependence of the $c_{ij}$ functions, nor of the constant $c$, on the control parameter $\beta$.
@@ 4 Maggiore GUP as an application
The mathematical relations we have derived in Section 3 are completely general and are based only on the request that the 2-form $\omega$ (2.4) is symplectic, without further assumptions on the $f$ and $L_{ij}$ functions.
Nevertheless, usually, in constructing a GUP theory other constraints are imposed on the general algebra (2). The most common additional demands include the rotational invariance of the algebra and the condition that the $L_{ij}$ functions satisfy an angular momentum algebra.
@@
$[\hat{q}_{i},\hat{q}_{j}]=\mp\frac{\hbar^{2}}{\kappa^{2}}i\epsilon_{ijk}\hat{J}_{k},$
$[\hat{q}_{i},\hat{p}_{j}]=i\hbar\delta_{ij}\sqrt{1\pm\tfrac{\hat{\rho}^{2}}{\kappa^{2}c^{2}}}.$
As it should be clear from the derivation, these specific GUP models acquire relevance since they are the most general models, defined only by the imposition of Jacobi identities and the inclusion of the spin operator. We stress that the generality of the model has to be intended in the sense specified above.
@@ 4.1 Closure condition and Maggiore algebra
$\begin{cases}(145)&\frac{\partial f}{\partial p_{2}}+a(p)\frac{\partial J_{3}}{\partial q_{1}}=0,\\ (146)&\frac{\partial f}{\partial p_{3}}-a(p)\frac{\partial J_{2}}{\partial q_{1}}=0,\\ (156)&\frac{\partial J_{1}}{\partial q_{1}}=0,\\ (245)&\frac{\partial f}{\partial p_{1}}-a(p)\frac{\partial J_{3}}{\partial q_{2}}=0,\\ (246)&\frac{\partial J_{2}}{\partial q_{2}}=0,\\ (256)&\frac{\partial f}{\partial p_{3}}+a(p)\frac{\partial J_{1}}{\partial q_{2}}=0,\\ (345)&\frac{\partial J_{3}}{\partial q_{3}}=0,\\ (346)&\frac{\partial f}{\partial p_{1}}+a(p)\frac{\partial J_{2}}{\partial q_{3}}=0,\\ (356)&\frac{\partial f}{\partial p_{2}}-a(p)\frac{\partial J_{1}}{\partial q_{3}}=0.\end{cases}$ (4.5)
@@
Accordingly, the gradient equation for $f$ that can be extracted from the system (4.5) becomes:
@@
By means of the expressions (4.1) for the $J$s functions we obtain a system of equations, which we can compactly write as:
@@
As a homogeneous linear system, its determinant is $f^{-2}(f^{2}+a(p_{1}^{2}+p_{2}^{2}+p_{3}^{2}))$. The singular points are represented by the equation:
$a(p)(p_{1}^{2}+p_{2}^{2}+p_{3}^{2})=-f^{2}(p).$ (4.15)
Clearly, this equation admits solution only for $a(p)<0$. If it is satisfied in an open subset of $\mathbb{R}^{3}$, then we can put the $a(p)$ derived from the above equation in the gradient equation (3.20). As a result, we obtain a unique solution for $f$, which is $f(p)=\kappa\sqrt{p_{1}^{2}+p_{2}^{2}+p_{3}^{2}}$. This solution is incompatible with the undeformed limit, since there is no any possible dependence of $\kappa$ on the deformation parameter $\beta$ which results in $f\xrightarrow[\beta\to 0]{}1$.
As a consequence, the equation (4.15) can be satisfied only on a submanifold with at least codimension one. Where in $\mathbb{R}^{3}$ the determinant is not singular, the unique solution of (4.14) is $\mathbf{s}=0$. Hence, since $s_{i}$ is a smooth function everywhere null but on a closed surface, for continuity it must be zero everywhere in $\mathbb{R}^{3}$.
@@
By employing the last obtained form of the $J$s we get: